On a 130 mile trip from Orlando to Sarasota, Roberto encountered a heavy thunderstorm for the last 40 miles of the trip. During the thunderstorm he averaged 20 miles per hour slower than before the storm. The entire trip took 2 1/2 hours. How fast did he travel before the storm?
2.5(20-x)=?x
50-2.5x=?x

Answer by mananth(16949)

total trip 130 miles
in storm 40 miles
..
ist leg 90 miles
..
first leg speed =x
second leg speed = x-20 mph
..
90/x + 40/(x-20)=2 1/2 hours
90/x + 40/(x-20)=5/2
LCD = x(x-20)
multiply by x(x-20)
90(x-20)+40x=5x(x-20)/2
multiply by 2
180(x-20)+80x=5x^2-100x
180x-3600+80x=5x^2-100x
5x^2-360x+3600=0
/5
x^2-72x+720=0
x^2-60x-12x+720=0
x(x-60)-12(x-60)=0
(x-60)(x-12)=0
x= 60 mph speed before the storm